OpenStudy (anonymous):

someone help please :( What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 6 x – y = 3 −1 5 −2 2

OpenStudy (anonymous):

There are several ways to solve this. The method I'd use is to double the bottom equation. If you double everything, the statement remains true. So you get 4x+2y=6 2x-2y=6 --------Now, add the two equations, so you get 6x=12 So x=2.

OpenStudy (anonymous):

i agree with vinnv226 the answer is 2 so therefore he is right

OpenStudy (anonymous):

thank you!

OpenStudy (jhannybean):

Solve one equation fory, lets say we take the second one. x - y = 3 Solving for y,we would subtract -x from both sides. -y = -x + 3 multiply the entire equation by -1 -1(-y = -x + 3) Simplify y = x - 3 Substitute into the first equation. 4x + 2y = 6 4x +2(x-3) = 6 Simplify 4x +2x -6 = 6 Add +6 to both sides 6x = 12 divide by 6 x = 2

OpenStudy (anonymous):

youre right! jhannybean! thank you

OpenStudy (jhannybean):

No problem :)

OpenStudy (anonymous):

can you help me with another equation?

OpenStudy (jhannybean):

Showed you two different methods of solving the same problem - elimination and substitution :D

OpenStudy (jhannybean):

Sure thing.

OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1150. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many one-step equations eliminated today? 430 720 1150 1300

OpenStudy (jhannybean):

Lets say the one step equations are x and the two step equations = y Your first statement: The number of one-step equations and two-step equations who have been eliminated today is equal to 115 We can rewrite this as the combination of both equations = total # :1150 \(\color{red}{x+ y = 1150}\) Second statement: If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300 \(\color{blue}{3x-2y=1300}\) Third statement : how many one-step equations eliminated today? \(\color{purple}{x=?}\)

OpenStudy (anonymous):

umm is it 720?

OpenStudy (jhannybean):

We can use @vinnv226 'smethod to evaluate this faster. x + y = 1150 3x - 2y = 1300 ------------- We want to eliminate the y's, so we multiply the whole equation by 2 2(x+y =1150) 3x - 2y = 1300 ------------------ 2x + 2y = 2300 3x - 2y = 1300 ------------------- 6x = 3600 x = 3600/6

OpenStudy (jhannybean):

your answer choices seem a bit off....

OpenStudy (anonymous):

I copied and pasted them exactly from the quiz

OpenStudy (anonymous):

@jhannybean you added 2x+3x and got 6x, you should have put 5x. So the answer is 3600/5 which is 720.

OpenStudy (anonymous):

thank you! :)

OpenStudy (jhannybean):

Oh yes you are right, whoops. then you are right, haha. Sorry about that.